# VPNeXt : Rethinking Dense Decoding for Plain Vision Transformer

### v2

@@ Abstract
We present VPNeXt, a new and simple model for the Plain Vision Transformer (ViT). Unlike the many related studies that share the same homogeneous paradigms, VPNeXt offers a fresh perspective on dense representation based on ViT. In more detail, the proposed VPNeXt addressed two concerns about the existing paradigm: (1) Is it necessary to use a complex Transformer Mask Decoder architecture to obtain good representations? (2) Does the Plain ViT really need to depend on the mock pyramid feature for upsampling? For (1), we investigated the potential underlying reasons that contributed to the effectiveness of the Transformer Decoder and introduced the Visual Context Replay (VCR) to achieve similar effects efficiently. For (2), we introduced the ViTUp module. This module fully utilizes the previously overlooked ViT real pyramid feature to achieve better upsampling results compared to the earlier mock pyramid feature. This represents the first instance of such functionality in the field of semantic segmentation for Plain ViT. We performed ablation studies on related modules to verify their effectiveness gradually. We conducted relevant comparative experiments and visualizations to show that VPNeXt achieved state-of-the-art performance with a simple and effective design. Moreover, the proposed VPNeXt significantly exceeded the long-established mIoU wall/barrier of the VOC2012 dataset, setting a new state-of-the-art by a large margin, which also stands as the largest improvement since 2015.
Index Terms:
Vision Transformer, Semantic Segmentation, Representation Learning
@@ I Introduction
First, research on the semantic segmentation decoder of plain ViT is heavily influenced by the original architectures of Transformer [36], DETR [37], and ViT [23], leading to high homogenization. In DETR, object detection is obtained through token query (cross-attention). In ViT, unlike the classification method used in the CNN era [38], it uses a class token to perform self-attention with other spatial tokens to obtain the classification. Many readers may have noticed that works on semantic segmentation decoders [24, 39, 40, 11, 12, 25] during the ViT era primarily focused on the DETR and ViT paradigms. Notable examples include Segmenter [24], MaX-DeepLab [39], and MaskFormer [11], whose decoders all share a similar name: Mask Decoder or Mask Transformer.
The strong effectiveness of the mask decoder does not need to be elaborated in this paper, as it has already demonstrated excellent feature regularization capabilities in these homogeneous works and has achieved very high mIOU on multiple semantic segmentation benchmark datasets [41, 42, 43].
@@
Besides, the plain ViT has an obvious disadvantage in semantic segmentation. Its tokenizer directly reduces the input image size by at least 16$\times$ times, which is not suitable for semantic segmentation tasks requiring the original size output. Mature solutions like FPN [44] and its variants [45, 46, 5, 47] cannot be implemented because they require multi-resolution pyramid features [48, 38, 49, 50], which plain ViT obviously cannot provide. To tackle the ViT’s resolution issue, a line of works [51] opt for transposed convolution or similar methods, which directly upsample feature maps in a stage-wise manner without reference to pyramid features. However, their impact is often limited as they still depend on the encoder to produce easily upsampling features, which is only slightly better than direct upsampling [24]. Another line of approaches involves creating a parallel pyramid network to produce high-resolution, low-level features while leveraging the pre-trained features of ViT. However, this results in a significant increase in inference overhead.
@@
To address these two questions, we propose VPNeXt (ViT context replay and upsample network; ’X’ represents new technology.), a completely new segmentation decoder that diverges from the popular paradigm without compromising performance.
Specifically, for the first question, VPNeXt includes a novel technology called Visual Context Replay (VCR) to achieve similar effectiveness as Mask-Decoder but with much greater efficiency, as VCR is only applied during training. VCR enables the same visual priors to be replayed during the early encoding stages of ViT. This allows for the interaction between fine visual priors and early features without increasing computational overhead during the inference stage. As a result, this approach leads to improved visual representations.
In response to the second question, we claim that plain ViT can also effectively extract native pyramid features, similar to those obtained by CNN networks [48, 38, 50] or Pyramid ViT networks [27, 52, 28], as opposed to the mock pseudo-pyramid features derived from resizing high-level features like SETR [51] and ViTDet [53]. Therefore, we present ViTUp, an effective technique that uncovers high-resolution features previously hidden in ViT and uses them to aid in final upsampling.
@@
In summary, the key contributions of this work are as follows:
  • We raised concerns about the inefficiency of the current heavy Mask-Decoder and the ineffectiveness of the upsampling paradigm for semantic segmentation in ViT.
@@
  • We proposed ViTUp to extract hidden high-resolution features in plain ViT to perform the pyramid-based upsampling.
  • The entire solution, VPNeXt, achieves state-of-the-art across multiple benchmark datasets with great computational efficiency.
  • VPNeXt also broke the long-standing mIoU wall of the VOC2012 dataset by a large margin, which also stands as the largest improvement since 2015.
@@ II-A Backbones for semantic segmentation
Due to the extremely high costs of pixel-level labeling, Current pre-trained models for natural scene images depend only on image-level labeling data, also known as the image classification backbone.
II-A1 Pre-training
Prior to 2020, most backbones [54, 48, 38, 49, 50, 55] utilized for downstream tasks were primarily pre-trained on the ImageNet-1K [56] dataset. Additionally, Google’s proposed backbones [50] were frequently pre-trained on their proprietary JFT-300M [57] dataset. After that, following the introduction of the Vision Transformer (ViT) [23] in late 2020, most modern backbone models [23, 27, 29, 58, 59] are now pre-trained on the ImageNet-21K dataset at a minimum. Some state-of-the-art backbones [60, 30, 61] have even been pre-trained on extensive datasets such as LAION-5B [62] or JFT-3B [63] (Google). These backbones are pre-trained on large image-level datasets, increasing diversity and enhancing downstream task generalization.
II-A2 Architecture
@@ IV Proposed Method
This work presents two methods. The first method is Visual Context Replay (VCR), which is a simple, efficient, and effective technique for enhancing the decoder’s input features. The second method is UpViT, which reveals the inherent high-resolution features that are typically thought to be absent in ViT [23].
@@ IV-A Visual Context Replay (VCR)
Figure 2: Our proposed VPNeXt consists of two main modules: VCR and ViTUp, which focus on enhancing features efficiently and addressing upsampling challenges respectively.
@@ VI-A Ablation studies on VCR
Mask decoder 2 (implicit) 67.88
w/o pyramid
@@
1007.62
TABLE IV: Comparisons to state-of-the-art methods on Pascal Context dataset. SS: Single-scale performance w/o flipping. MF: Multi-scale performance w/ flipping. “-” in column SS indicates that this result was not reported in the original paper.
@@
IDRNet [78] Swin-L NIPS’23 - 64.5
@@
InternImage [81] InternImage-H CVPR’23 - 70.3
@@ VI-D Compare with state-of-the-arts
TABLE V: Comparisons to state-of-the-art methods on COCOStuff164k dataset. SS: Single-scale performance w/o flipping. MF: Multi-scale performance w/ flipping. “-” in column SS or MF indicates that this result was not reported in the original paper.
@@ VIII Experiments on Cityscapes Dataset
TABLE VI: Comparisons to state-of-the-art methods on Cityscapes validation set. SS: Single scale performance w/o flipping. MF: Multi-scale performance w/ flipping. “-” in column SS indicates that this result was not reported in the original paper.
@@
SETR [52] ViT-L CVPR’21 - 82.2
Segmenter [24] ViT-L ICCV’21 - 81.3
OCR [9, 85] HRFormer-B NIPS’21 - 82.6
@@ IX Experiments on VOC2012
[…]
@@ X Visualization
Figure 3: Visual comparison between Mask2Former + InternImage-H [12, 81] and our proposed VPNeXt on the COCOStuff164k [42] dataset shows that VPNeXt achieves superior segmentation results, particularly in challenging categories such as food.
Figure 4: Visualization analysis of the intermediate feature map for the VCR is based on the 15th layer of the ViT. At the pixel position marked by the red dot, the replay-optimized feature map displays significantly stronger and more detailed semantic information concerning intra-class pixels.
X-A Visualization on COCOStuff164k dataset
We first conducted a visualization comparison on the challenging COCOStuff164k dataset [42]. As shown in Figure. 3, our proposed VPNeXt achieves significantly better segmentation results compared to the state-of-the-art Mask2Former + InternImage-H [12, 81], particularly in some challenging categories, such as food.
X-B Visualization on intermediate feature map
We conducted a visualization analysis of the intermediate feature map of the VCR. For this analysis, we utilized the feature map from the $15^{\text{th}}$ layer of the ViT. As illustrated in Figure. 4, at the position marked by the red dot ($\bullet$), the replay-optimized feature map presents significantly stronger and more intensive semantic information regarding intra-class pixels.

### v3

@@ Abstract
We present VPNeXt, a new and simple model for the Plain Vision Transformer (ViT). Unlike the many related studies that share the same homogeneous paradigms, VPNeXt offers a fresh perspective on dense representation based on ViT. In more detail, the proposed VPNeXt addressed two concerns about the existing paradigm: (1) Is it necessary to use a complex Transformer Mask Decoder architecture to obtain good representations? (2) Does the Plain ViT really need to depend on the mock pyramid feature for upsampling? For (1), we investigated the potential underlying reasons that contributed to the effectiveness of the Transformer Decoder and introduced the Visual Context Replay (VCR) to achieve similar effects efficiently. For (2), we introduced the ViTUp module. This module fully utilizes the previously overlooked ViT real pyramid feature to achieve better upsampling results compared to the earlier mock pyramid feature. This represents the first instance of such functionality in the field of semantic segmentation for Plain ViT. We performed ablation studies on related modules to verify their effectiveness gradually. We conducted relevant comparative experiments and visualizations to show that VPNeXt achieved state-of-the-art performance with a simple and effective design. Moreover, the proposed VPNeXt significantly exceeded the long-established mIoU wall/barrier of the VOC2012 dataset, setting a new state-of-the-art by a large margin, which also stands as the largest improvement since 2015. The code will be available on https://github.com/edwardyehuang/research/VPNeXt.
@@ I Introduction
First, many works on semantic segmentation decoders [24, 36, 37, 11, 12, 25] during the ViT era primarily focused on the DETR [38] and ViT paradigms. Notable examples include Segmenter [24], MaX-DeepLab [36], and MaskFormer [11], whose decoders all share a similar name: Mask Decoder or Mask Transformer. The strong effectiveness of the mask decoder does not need to be elaborated in this paper, as it has already demonstrated excellent feature regularization capabilities by achieving very high mIOU on multiple benchmark datasets [39, 40, 41].
@@
[…]
@@
To address these two questions, we propose VPNeXt (ViT context replay and upsample network; ’X’ represents new technology.). Specifically, for the first question, VPNeXt includes a novel technology called Visual Context Replay (VCR) to achieve similar effectiveness as Mask-Decoder but with much greater efficiency, as VCR is only applied during training. VCR enables the same visual priors to be replayed during the early encoding stages of ViT. This allows for the interaction between fine visual priors and early features without increasing computational overhead during the inference stage. As a result, this approach leads to improved visual representations.
In response to the second question, we claim that plain ViT can also effectively extract native pyramid features, similar to those obtained by pyramid networks [46, 47, 49, 27, 51, 28], as opposed to the mock pseudo-pyramid features derived from resizing high-level features like SETR [50] and ViTDet [52]. Therefore, we present ViTUp, an effective technique that uncovers high-resolution features previously hidden in ViT and uses them to aid in final upsampling.
@@
In summary, our key contributions include :
@@
  • We proposed ViTUp to extract native high-resolution features in plain ViT for the pyramid-based upsampling.
  • The entire solution, VPNeXt, achieves state-of-the-art across multiple benchmark datasets, and also broke the long-standing mIoU wall of the VOC2012 dataset by a large margin, which also stands as the largest improvement since 2015.
@@ IV Proposed Method
This work presents two methods. The first is Visual Context Replay (VCR), which is a simple, efficient, and effective technique for enhancing the decoder’s input features. The second is UpViT, which reveals the inherent high-resolution features that are typically thought to be absent in ViT [23].
@@ IV-A Visual Context Replay (VCR)
Figure 2: Our proposed VPNeXt consists of two main modules: VCR and ViTUp, which focus on enhancing features efficiently and addressing upsampling challenges, respectively.
@@ VI-A Ablation studies on VCR
Mask decoder w/o pyramid 2 (implicit) 67.88
@@
552.85
TABLE IV: Comparisons to state-of-the-art methods on Pascal Context dataset. $S$: Single-scale performance w/o flipping. MF: Multi-scale performance w/ flipping. “-” in column $S$ indicates that this result was not reported in the original paper.
@@
IDRNet [69] Swin-L NIPS’23 63.8 64.5
@@
InternImage [72] InternImage-H CVPR’23 69.7 70.3
SegViTv2 [73] ViT-L IJCV’23 - 67.1
@@ VI-D Compare with state-of-the-arts
TABLE V: Comparisons to state-of-the-art methods on COCOStuff164k dataset. $S$: Single-scale performance w/o flipping. MF: Multi-scale performance w/ flipping.
@@ VIII Experiments on Cityscapes Dataset
TABLE VI: Comparisons to state-of-the-art methods on Cityscapes validation set. $S$: Single scale performance w/o flipping. MF: Multi-scale performance w/ flipping.
@@
SETR [51] ViT-L CVPR’21 79.2 81.0
Segmenter [24] ViT-L ICCV’21 79.1 81.3
OCR [9, 78] HRFormer-B NIPS’21 81.9 82.6
@@ IX Experiments on VOC2012
It is worth noting that the training strategies employed, including data augmentation techniques and the use of the combined VOC2012 [35] and SBD [91] datasets (also known as ‘trainaug’), are consistent with those used in existing works, such as DeepLabV3+ [92].
TABLE VII: New breakthroughs in the VOC2012 leaderboard! Due to limited space on the page, we have simplified some category names (e.g., "Aero Plane" to "Plane") and only listed the top 15 methods. Zoom in to see better. To view the full leaderboard, please visit http://host.robots.ox.ac.uk:8080/leaderboard/displaylb_main.php?challengeid=11&compid=6. (The main website might be down, see web archive mirror: https://web.archive.org/web/20250620175614/http://host.robots.ox.ac.uk:8080/leaderboard/displaylb_main.php?challengeid=11&compid=6)
@@ X Visualization
Figure 3: Visualization analysis of the intermediate feature map for the VCR, based on the 15th layer of the ViT on the Pascal Context dataset, at the pixel position marked by the red dot, the replay-optimized feature map displays significantly stronger and more detailed semantic information concerning intra-class pixels. Note that, in order to maintain controlled variables, modules other than VCR, such as UpViT are not utilized.
Figure 4: Visual comparison between Mask2Former + InternImage-H [12, 72] and our proposed VPNeXt on the COCOStuff164k [40] dataset. VPNeXt achieves superior segmentation results, particularly in challenging categories.
Figure 5: Visual comparisons between SegNeXt, Segmenter and our proposed VPNeXt on the Cityscapes val set. Zoom in to see better. We highlight the noticeable shortcomings of the SegNeXt and Segmenter approaches.
X-A Visualization on intermediate feature map
We first conducted a visualization analysis of the intermediate feature map of the VCR. For this analysis, we utilized the feature map from the $15^{\text{th}}$ layer of the ViT. As illustrated in Fig. 3, at the position marked by the red dot ($\bullet$), the replay-optimized feature map presents significantly stronger and more intensive semantic information regarding intra-class pixels.
X-B Visualization on COCOStuff164k dataset
We then conducted a visualization comparison on the challenging COCOStuff164k dataset [40]. As shown in Fig. 4, our proposed VPNeXt achieves significantly better segmentation results compared to the state-of-the-art Mask2Former + InternImage-H [12, 72], particularly in some challenging categories, such as food.
X-C Visualization on Cityscapes dataset
Finally, we provide qualitative visual comparisons among our proposed VPNeXt, SegNeXt, and Segmenter. As illustrated in Fig. 5, our proposed VPNeXt successfully addresses the upsampling issues of the native ViT and even produces finer details than the pyramid-based SegNeXt.